OpenStudy (anonymous):

In a foreign country, half of 5 = 3. Based on that same proportion, what's one-third of 10

OpenStudy (anonymous):

In a foreign country, half of 5 = 3. That means 2.5 = 3. 2.5 : 3 is our ratio / proportion. 1/3 of 10 means 3.33. That means we have to solve for x. 2.5 : 3 3.33 : x 2.5x = 3.33*3 2.5x = 10 x = 4

OpenStudy (anonymous):

wrong it is 2. gotcha.

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

2. (or maybe 288 if life is horrible) So, 48/2(9+3) = 48/2(12)=48/24=2. Unless you go 48/2(9+3) = 48/2(12)=24(12)=288, in which case, life IS horrible. I did fall into the rabbit hole.... this question is mean. mean but enlightening!

OpenStudy (anonymous):

accept defeat from a 13 year old, @kevinzhengli ?